OpenStudy (anonymous):

simplify the inverse trig function with answers (-inf, +inf): cos(inverse)(sqrt3/2)

OpenStudy (anonymous):

answers from \[(-\infty,\infty)\] equation is \[\cos^{-1} (\sqrt{3}/2)\]

OpenStudy (anonymous):

Cosine takes in angles and gives out a ratio of sides. Inverse cosine goes backwards. It takes in a ratio of sides, and it tells you what angles have that ratio.

OpenStudy (anonymous):

And cos is a/h, so just make a triangle with those sides, then figure out what the angle is.

OpenStudy (anonymous):

still have your unit circle or you got it memorized?

OpenStudy (anonymous):

I wouldn't use the unit circle here, since the hypotenuse is 2.

OpenStudy (anonymous):

But I suppose you could.

OpenStudy (anonymous):

sqrt3/2 is a cos/sin of one of the common angles though right?

OpenStudy (anonymous):

probably... maybe. Lols.

OpenStudy (anonymous):

i have the unit circle i know most of it by heart

OpenStudy (anonymous):

ok... cos(what angle) = sqrt3/2 ???

OpenStudy (anonymous):

wait cos is the x-value right? o.o

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

30 degrees?

OpenStudy (anonymous):

yes... but what else?

OpenStudy (anonymous):

300 degrees?

OpenStudy (anonymous):